OpenStudy (anonymous):

Given the transformation u=y-lnx and v=y-x+1 ... Calculate the partial derivatives dy/dv and dx/du

OpenStudy (anonymous):

If you mean \(\frac{\partial y}{\partial v}\) and \(\frac{\partial x}{\partial u}\), then you need to solve the system of equations for explicit equations for \(y\) and \(x\) in terms of \(u\) and \(v\), and then just find the partials as you always would.

OpenStudy (anonymous):

can you write the equations please ? i don't really get it ..

hartnn (hartnn):

since u are finding partial derivative of v w.r.t y u treat x as constant in v=y-x+1 so now when u diff w.r.t y what u get??

OpenStudy (anonymous):

i have du = dy -dx(1/x) and dv = dy - dx ... but how do i proceed ?

hartnn (hartnn):

but i told u x is constant in v=y-x+1 so dv/dy is just 1

OpenStudy (anonymous):

alright .. i get it so i have dv/dy =1 and how do i get dx/du ?

hartnn (hartnn):

treat y as constant in u=y-ln x

OpenStudy (anonymous):

Thanks ! but how do you know that you should treat x and y as constants each time ?

hartnn (hartnn):

because of the word 'partial' the only variable would be the one u are diff with.

OpenStudy (anonymous):

!!! thank you !!

hartnn (hartnn):

welcome :)